An archer shoots an arrow straight up at 76.7 m/s. (a) How high does the arrow rise? (b) How much time is the arrow in the air? (c) What is the arrow's velocity when it reaches the same height that it was shot from?
Answer by Alan3354(69443)

An archer shoots an arrow straight up at 76.7 m/s.
(a) How high does the arrow rise?
You didn't spec what value to use for gravity.
Using 10 m/sec/sec:
The rise time = 76.7/10 = 7.67 seconds
Its average speed is 76.7/2 = 38.35 m/sec
max height = 7.67*38.35 = 294.1445 meters
============================================
(b) How much time is the arrow in the air?
Rise and fall time are equal.
7.67*2 = 15.34 seconds
------------------------
(c) What is the arrow's velocity when it reaches the same height that it was shot from?
Same as the launch speed = 76.7 m/sec
Speed, not velocity.